OpenStudy (anonymous):

write the function that models each variation. find z when x=4 and y=9. z varies directly with x and inversely with y. when x=6 and y=2, z=15.

OpenStudy (anonymous):

z varies directly with x and inversely with y. z=kx/y, where k is any constant x=6, y=2, z=15 15 = k(6)/2 15 = 3k k=5 z=3x/y when x=4 & y=9 z= 3(4)/9 = 12/9 = 4/3

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

yes.thanks

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

yeah